OpenStudy (yrelhan4):

how is log[x+ {(1+x^2)^1/2}] an odd function?

OpenStudy (anonymous):

seems unlikely to me

OpenStudy (anonymous):

for example \(f(1)=1+\sqrt{2}\) and \(f(-1)=-1+\sqrt{2}\) so it is not odd

OpenStudy (yrelhan4):

i am sorry. just corrected the question.

OpenStudy (anonymous):

ok now it is odd

OpenStudy (yrelhan4):

@satellite73 how? i put -x in place of x? log[-x+ {(1+x^2)^1/2}]. thats what i get.

OpenStudy (anonymous):

[-x+ {(1+x^2)^1/2}]*[x+ {(1+x^2)^1/2}]=1 hope it is enough hint

OpenStudy (anonymous):

yeah i did that too and it didn't help maybe multiply rationalize the numerator inside the log

OpenStudy (anonymous):

hey mann this is done

OpenStudy (anonymous):

i like @RajshikharGupta solution more than mine

OpenStudy (anonymous):

thank u

OpenStudy (yrelhan4):

thank you!

OpenStudy (anonymous):

2*thank u